# CoT-Space: A Theoretical Framework for Internal Slow-Thinking via Reinforcement Learning

### v1

@@ 2.3 Formulation of CoT-Space
(CoT reasoning state.) Given a query $q$ and an incomplete set of intermediate reasoning steps $\xi_{o}$, the current reasoning state is defined as $s_{o}=(q,\xi_{o})$.
Definition 2.7 establishes a reasoning-level state that is a more meaningful unit of analysis than a single token. We then introduce a special class of states that represent successful problem solutions, which we call minimums.
Definition 2.8.
(Minimums in reasoning space.) Given a query $q$ with a definite golden answer $\phi$. Assume $\Xi_{q}$ is the set of all reasonable intermediate reasoning process for query $q$. Then the minimums of the reasoning space are defined as $\bm{M}_{q}=\left\{(q,\bm{\xi},\phi)~|~\bm{\xi}\in\Xi_{q}\right\}$.
According to Definition 2.8, minimums are special states where the query has been successfully solved. For any intermediate reasoning state, we can then define which minimums are reachable from it.
Definition 2.9.
(Reachable minimums.) Given a query $q$ with a definite golden answer $\phi$, and an incomplete set of intermediate reasoning steps $\xi_{o}=[\xi_{1},\xi_{2},\dots,\xi_{t}]$. The reachable minimums are defined as $\bm{R}^{o}_{q}=\left\{(q,\bm{\xi},\phi)~|~\xi_{o}\sqsubseteq\bm{\xi}~\text{and}~\bm{\xi}\in\Xi_{q}\right\}$, where $a\sqsubseteq b$ represents that $a$ is prefix of $b$.
Definition 2.10.
(Reachable distance and nearest reachable minimum.) Given a query $q$ with a definite golden answer $\phi$, and a reasoning state $s_{o}$ on this query, the reachable distance between the $s_{o}$ and a reachable minimum is defined as $\operatorname{dist}(s_{o},m)=|\bm{\xi}|-|\xi_{o}|$, where $m\in\bm{R}^{o}_{q}$, $\bm{\xi}$ and $\xi_{o}$ are intermediate steps of $m$ and $s_{o}$ respectively. The nearest minimum is defined as $m^{*}_{o}=\underset{m\in\bm{R}^{o}_{q}}{\arg\min}\,\operatorname{dist}(s_{o},m)$.
Definition 2.11.
(Reasoning loss.) Given a query $q$ with a definite golden answer $\phi$, and a reasoning state $s_{o}$ on this query, there exists a measure $C$ that quantifies the loss of $s_{o}$ towards the golden answer, which satisfies: for two different states $s_{i},s_{j}$, $C(s_{i})<C(s_{j})$ if and only if $\operatorname{dist}(s_{i},m^{*}_{i})<\operatorname{dist}(s_{j},m^{*}_{j})$. Specially, $\forall m\in\bm{M}_{q},C(m)=0$.
Definition 2.11 conceptualizes the reasoning process as a descent in a loss landscape. The function $C(\cdot)$ acts as a proxy for the distance to a correct solution, where a lower loss indicates a more advanced stage of reasoning. For instance, consider the problem in Figure 1. A state $s_{i}$ with the intermediate thought “He writes $12$ pages every week” is closer to the final answer than a state $s_{j}$ with only “He writes each friend $32=6$ pages a week.” According to this definition, because $s_{i}$ requires fewer steps to reach a minimum (i.e., $\operatorname{dist}(s_{i},m_{i}^{*})<\operatorname{dist}(s_{j},m_{j}^{*})$), its reasoning loss must be lower, satisfying $C(s_{i})<C(s_{j})$.
With this complete reasoning-level perspective, we can now model the LLM reasoning process as an optimization problem rather than a token-level MDP. This formulation simplifies the analytical foundation by reframing the process as an update to the reasoning state, guiding it toward one of the solution minima.
@@ 3 Analytical Applications for Overthinking within CoT-Space
This section demonstrates the analytical utility of the CoT-Space framework by applying it to a critical issue in LLM reasoning: the existence of an optimal thought process length and the related problem of “overthinking”. Moving beyond traditional analyses, we utilize the continuous perspective of CoT-Space to provide a theoretical foundation for this empirical phenomenon. Our analysis proceeds in two parts: we first examine the reasoning process through the lens of optimization and noise (Subsection 3.1), and second, we assess its generalization capabilities from the perspective of risk (Subsection 3.2). Finally, we briefly conclude our theoretical insights in Subsection 3.3. This application of CoT-Space showcases its potential to ground empirical findings in a robust theoretical model.
@@ 3.1 Reasoning as Optimization: A Perspective of Noise
Figure 3: An illustration of LLM reasoning modeled as an optimization process. (a) In traditional ML, the goal is parameter optimization, where an algorithm iteratively calculates a gradient from data to update model parameters and converge to a minimum in the loss landscape. (b) In the CoT-Space framework, LLM reasoning is framed as state optimization, where the policy takes the current state and instruction to generate a new reasoning step, thereby updating the state to move closer to a solution minimum. (c) This panel visualizes the optimization process on a reasoning loss landscape, analogizing the CoT length to the learning rate in ML. A suitable CoT length, much like an optimal learning rate, navigates the space effectively to converge to a high-quality, generalizable solution.
By treating reasoning as an optimization process, we can draw a direct comparison to traditional machine learning (ML), as illustrated in Figure 3. As shown in subgraph (a), traditional ML optimization algorithms, such as SGD or Adam, formulate a cycle where the algorithm calculates a gradient based on the current parameters and training data to iteratively update the model. Similarly, in subgraph (b), when an LLM performs reasoning, its internal policy takes the current reasoning state and query instruction as input. It then generates a new step, which updates the state by concatenating the new output. This process establishes a parallel optimization cycle for LLM reasoning.
Such a comparison inspires an application of the sharpness-aware minimization theorem from classical ML. This theorem posits that an optimal amount of noise in the optimization process leads to a converged minimum with better generalization ability. As shown in subgraph (c) with the blue legends, the learning rate can introduce noise into the gradient descent process. A learning rate that is too large will fail to converge, while one that is too small may converge to a sharp, low-quality minimum. Only a suitable learning rate introduces the right amount of noise to converge to a flat, high-quality minimum. Given the similarities in the nature of optimization, this theoretical analysis can be applied to the reasoning process of LLMs, as shown with the orange legends in subgraph (c). In this analogy, the CoT length is analogous to the learning rate. A CoT that is too short may fail to produce a correct answer, while a CoT that is too long may lead to “overthinking”, resulting in a lower-quality response. Therefore, a suitable CoT length is essential for generating a high-quality response.

### v2

@@ 2.3 On the Continuum Nature of the Reasoning-Level State Space
[…]
We will further show the convergence relationship between the continuum errors and the density, thereby revealing the rationality of the reasoning-level state space continuous approximation.
Theorem 2.12.
(Convergence of the continuum error). The expected angular error $\mathbb{E}\left[\epsilon_{A}(s)\right]$ and expected magnitude error $\mathbb{E}\left[\epsilon_{M}(s)\right]$ are both upper-bounded by a function that is inversely related to the reasoning state density $\rho(K)$: $\mathbb{E}[\epsilon_{A}(s)]\leq\mathcal{O}\left(\rho(K)^{-\frac{1}{D}}\right)$ and $\mathbb{E}[\epsilon_{M}(s)]\leq\mathcal{O}\left(\rho(K)^{-\frac{1}{D}}\right).$
[…]
With the continuum nature of the reasoning space established, we bridge the gap between discrete language generation and classical learning theories. This continuum approximation is not merely a conceptual shift, it is the critical step that licenses the use of powerful analytical tools. Specifically, it allows us to conceptualize a smooth reasoning loss landscape and thus define a corresponding gradient, enabling us to model the generation of a new reasoning step as a descent-like update.
Building upon this, we can frame the entire reasoning process as a trajectory in a continuous state space, and link many LLM phenomena, such as hallucination, prompt sensitivity, emergent abilities, and effectiveness of external slow-thinking strategies, with classical theoretical results. A brief discussion is provided in Appendix E. In the next section, we take the “overthinking” as a detailed example to reveal the potential of CoT-Space for revitalizing the classical theoretical results.
@@ 3 Analytical Applications for Overthinking within CoT-Space
This section utilizes the CoT-Space framework and combines classic learning theories to analyze the “overthinking” phenomenon, thus yielding the potential of CoT-Space. We provide a theoretical foundation by examining the issue through two lenses: optimization noise (Subsection 3.1) and generalization risk (Subsection 3.2), followed by a summary of our theoretical insights (Subsection 3.3).
@@ 3.1 Reasoning as Optimization: A Perspective of Noise
Figure 4: Modeling LLM reasoning as an optimization process. (a) Traditional ML performs parameter optimization by updating model weights. (b) Our framework models reasoning as state optimization, where new reasoning steps iteratively update the current state. (c) This panel visualizes this process on a reasoning loss landscape, where the CoT length is analogous to a learning rate; an optimal value is crucial for converging to a high-quality, generalizable solution.
Leveraging the continuous nature of CoT-Space, we frame LLM reasoning as an optimization process analogous to ML, as shown in Figure 4. In this analogy, an LLM policy iteratively refines its reasoning state towards a solution, much like an optimizer updates parameters. This perspective allows us to apply classical theories like sharpness-aware minimization, which suggests that an optimal amount of noise promotes convergence to better solutions. The CoT length functions as a learning rate that controls optimization noise: if too short, the process fails to converge; if too long, it leads to “overthinking”, akin to finding a sharp, poor-quality minimum.
Building on this, we can model the optimization of LLM reasoning as a form of reasoning loss descent and perform a theoretical analysis of the relationship between the noise scale and CoT length.
Theorem 3.1.
The proof is provided in Appendix B.4. Theorem 3.1 clarifies the relationship between the noise $g$ and the CoT length $L$ in the same reasoning process. According to the sharpness-aware minimization theorems, given a reasoning task, there exists an optimal noise scale $g_{\text{opt}}$ in the reasoning process, which corresponds to an optimal CoT length $L_{\text{opt}}\propto\frac{1}{g_{\text{opt}}}$. In essence, this theoretical analysis provides a fundamental basis for why an optimal CoT length is crucial for effective reasoning, highlighting its role in introducing the right amount of noise to achieve a flat, high-quality, and generalizable solution.
3.2 Reasoning and Generalization: A Perspective of Risk
The previous section considered reasoning from the perspective of noise. Here, we analyze the generalization risk of a policy to provide further theoretical insights, focusing on the upper bounds of reasoning risks. In the context of CoT-Space, risks can be defined similarly to classical ML. The total expected error of a policy $\pi$ on the true data distribution $\mathcal{D}$, denoted by $R(\pi)=\mathbb{E}_{q\sim\mathcal{D}}[C(s_{q}^{\pi})]$, can be decomposed into two distinct components: $R(\pi)=\hat{R}_{n}(\pi)+\left(R(\pi)-\hat{R}_{n}(\pi)\right)$. Here, $\hat{R}_{n}(\pi)=\frac{1}{n}\sum_{i=1}^{n}C(s_{q_{i}}^{\pi})$ is the empirical loss (a.k.a. bias) on the training set. $R(\pi)-\hat{R}_{n}(\pi)=\mathbb{E}_{q\sim\mathcal{D}}[C(s_{q}^{\pi})]-\frac{1}{n}\sum_{i=1}^{n}C(s_{q_{i}}^{\pi})$ is the reasoning generalization error (a.k.a. virance). In the following parts, we will analyze how each component behaves as a function of the CoT complexity.
3.2.1 An Upper Bound for Generalization Error: The Overfitting Risk
Given $n$ queries $\{q_{i}\}_{i=1}^{n}$ sampled i.i.d. from a distribution $\mathcal{D}$, and a policy $\pi$ learned from $\{q_{i}\}_{i=1}^{n}$, let’s denote $s_{q}^{\pi}=\pi(q)$ as the end reasoning state output by $\pi$ on query $q$. The reasoning generalization error of $\pi$ on $\mathcal{D}$ is defined as: $R(\pi)-\hat{R}_{n}(\pi)=\mathbb{E}_{q\sim\mathcal{D}}[C(s_{q}^{\pi})]-\frac{1}{n}\sum_{i=1}^{n}C(s_{q_{i}}^{\pi}).$ This error quantifies a policy’s ability to generalize its reasoning to unseen queries, a concept parallel to the generalization error in traditional machine learning theory. By applying classical PAC-Bayesian and information-theoretic results of ML, we can derive an upper bound for the reasoning generalization error.
Theorem 3.2.
A detailed discussion and proof are provided in Appendix B.5. Theorem 3.2 explicitly connects the generalization error to the observable, structural properties of the reasoning process. It suggests that overfitting can occur in overlong response length. This provides a direct theoretical basis for regularizing the CoT generation process in RL to improve reasoning generalization.
3.2.2 A Lower Bound for Empirical Loss: The Underfitting Risk
Building upon this intuition, we can derive a lower bound for the empirical reasoning loss $\hat{R}_{n}(\pi)$.
Theorem 3.3.
$\hat{R}_{n}(\pi)\geq P_{\pi}(L<L^{*})\cdot C_{fail},$
where $L^{*}$ is the minimum required CoT depth for successful reasoning, $P_{\pi}(L<L^{*})=n_{fail}/n$ is the empirical failure rate of policy $\pi$ on the training set. $C_{fail}$ is the upper bound of the reasoning loss in the failed reasoning results.
A detailed proof is provided in Appendix B.6. Theorem 3.3 provides a formal basis for the “underfitting” portion of our trade-off curve. It demonstrates that the empirical loss is directly constrained by the probability that the policy generates a CoT that is too short for the problem’s intrinsic complexity. A policy that tends to produce short CoTs will have a high failure rate on non-trivial problems. Consequently, its empirical loss $\hat{R}_{n}(\pi)$ is guaranteed to be high. This formalizes why a policy must maintain enough CoT length $L$ to achieve a low empirical loss, thus completing our explanation for the tradeoff in our total error analysis and the existence of an optimal $L_{\text{opt}}$.
3.3 Takeaways
Our previous analysis from the perspectives of noise (Subsection 3.1) and risk (Subsection 3.2) yields the following key conclusions about the factors governing the convergence of the optimal CoT length, $L_{\text{opt}}$, in LLM reasoning via RL.
Remark 1. The intrinsic difficulty of a task dictates $L_{\text{opt}}$. Per Theorem 3.3, more complex tasks require a greater reasoning depth ($L^{*}$) to avoid underfitting. Consequently, an effective policy must learn a sufficiently long $L_{\text{opt}}$ commensurate with the task’s complexity to minimize empirical loss.
Remark 2. Model capacity is inversely related to the optimal CoT length due to overfitting risk. According to Theorem 3.2, higher-capacity models face a greater generalization error, which scales with CoT length $L$. To mitigate this risk, more powerful models learn to produce more concise reasoning, converging to a shorter $L_{\text{opt}}$ to regularize their outputs.
Remark 3. The converged CoT length is independent of the specific RL algorithm used. Our analysis indicates that $L_{\text{opt}}$ is an intrinsic property determined by task difficulty and model capacity. The algorithm’s role is more likely to effectively optimize the policy to find this good solution.
Remark 4. The optimization noise scale is inversely correlated with the converged CoT length. Theorem 3.1 shows that higher noise levels incentivize convergence to a shorter $L_{\text{opt}}$, as noisier environments favor more robust, concise reasoning paths that are less susceptible to disruption.
To validate these theoretical claims, the following section presents a series of experiments designed to systematically test each of these four remarks.
4 Experiments
This section presents a series of experiments designed to empirically validate the theoretical insights established in the preceding sections. We systematically investigate the relationship between the optimal CoT length ($L_{\text{opt}}$) and key factors. The detailed experimental setup is listed in Appendix C
Figure 5: Empirical validation of factors governing optimal CoT length ($L_{\text{opt}}$). (a) Task difficulty: $L_{\text{opt}}$ increases with task complexity. (b) Model capacity: More powerful models converge to a shorter $L_{\text{opt}}$ to mitigate overfitting. (c) RL algorithm: The converged $L_{\text{opt}}$ is largely algorithm-agnostic. (d) Optimization noise: $L_{\text{opt}}$ is inversely correlated with noise (proxied by KL-Loss).
Task Difficulty and $L_{\text{opt}}$. To validate Remark 1, we test on tasks of increasing difficulty (GSM8K and five levels of MATH) with Qwen2.5-7B-Base. As shown in Figure 5(a), $L_{\text{opt}}$ demonstrates a clear monotonic increase with task complexity. This aligns with Theorem 3.3, as more challenging tasks require a greater reasoning depth ($L^{*}$). To avoid the high empirical loss from underfitting, the policy learns to generate a correspondingly longer CoT. Additional validations are in Appendix D.
Model Capacity and $L_{\text{opt}}$. To verify Remark 2, we examine models with varying capacities. Figure 5(b) reveals a distinct negative correlation: within each model family (e.g., Llama 3, Qwen 2.5, and Qwen 3), larger and higher-performing models consistently converge to a shorter $L_{\text{opt}}$. This finding supports our analysis based on Theorem 3.2. Larger models are more prone to overfitting, a risk exacerbated by longer CoTs. Consequently, they learn to produce more concise reasoning to minimize total error by balancing empirical performance with generalization.
RL Algorithm and $L_{\text{opt}}$. To test Remark 3, we compare five different RL algorithms. As shown in Figure 5(c), despite different learning dynamics, all algorithms guide the policy to a similar final $L_{\text{opt}}$ for a given task and model, even though the initial dynamics differ. This supports our claim that the RL algorithm primarily facilitates optimization, while the final $L_{\text{opt}}$ is an intrinsic property determined by task difficulty and model capacity, not the specific algorithm used.
Noise Scale and $L_{\text{opt}}$. Finally, to validate Remark 4, we use the KL-Loss as a proxy for optimization noise. Figure 5(d) shows a clear inverse correlation between the final KL-Loss and $L_{\text{opt}}$. This result substantiates Theorem 3.1, confirming that noisier optimization environments favor shorter CoTs to ensure robust convergence, as longer reasoning paths are more susceptible to disruption.
5 Related Work
Reasoning with LLMs. Prompting techniques have advanced LLM reasoning beyond few-shot learning (Brown et al., 2020). Chain-of-Thought (CoT) prompting elicits intermediate reasoning steps (Wei et al., 2022b), with methods like Self-Consistency improving robustness (Wang et al., 2022). Other reasoning formats include Program-of-Thought (Chen et al., 2022) and internal thought tokens (Zelikman et al., 2024; Chen et al., 2025a). While theoretical analyses are emerging (Ton et al., 2024), much of this work remains empirical and lacks a unifying theoretical foundation.
RL for LLM Reasoning. Reinforcement Learning (RL) is a key method for test-time scaling, a technique proven to enhance LLM reasoning (Snell et al., 2024; Wu et al., 2024; Chen et al., 2025b). Research in this area follows two main tracks: developing new post-training frameworks like VC-PPO, DAPO, and VAPO (Yuan et al., 2025; Yu et al., 2025; Yue et al., 2025b), and understanding the fundamental effects of RL. Studies in the latter track investigate RL’s advantages over SFT (Setlur et al., 2025), its impact on knowledge boundaries (Yue et al., 2025a), the role of reward modeling (Shao et al., 2025), and its tendency to amplify pre-trained behaviors (Zhao et al., 2025). The absence of a common theoretical framework, however, can lead to divergent findings.
Optimization and Generalization. A core concept in machine learning optimization is that flatter minima in the loss landscape often lead to better generalization, a principle central to Sharpness-Aware Minimization (SAM) (Keskar et al., 2016). This effect is closely linked to the noise inherent in the optimization process (Smith & Le, 2018). Generalization itself, which measures performance on unseen data, is often analyzed through the powerful lens of information theory. Concepts like entropy and mutual information are used to establish theoretical bounds on the generalization capacity of deep learning models (Russo & Zou, 2019; Xu & Raginsky, 2017) and have recently been applied to analyze synthetic data generation and reasoning errors in LLMs (Gan & Liu, 2024; Ton et al., 2024).
6 Conclusion
In this paper, we introduced CoT-Space, a novel theoretical framework that addresses the critical misalignment between token-level RL theories and the reasoning-level nature of CoT by reframing LLM reasoning as a continuous optimization problem. This shift in perspective serves as a conceptual bridge, demonstrating that the foundational principles of classical machine learning are not obsolete but can be powerfully revitalized to analyze modern LLM behaviors. By revitalizing classical concepts of optimization and risk analysis, we demonstrate that a fundamental trade-off between underfitting (shorter CoTs) and overfitting (longer CoTs) governs the reasoning process. This analysis provides a solid theoretical grounding for the existence of an optimal CoT length, $L_{opt}$, that minimizes total error. Looking forward, the CoT-Space framework offers a foundation for designing more principled algorithms and advancing our ability to build more capable and reliable AI reasoning systems.
Appendix A Use of Large Language Models
In line with the ICLR policy, we disclose that Large Language Models (LLMs) were used as a general-purpose writing assistant during the preparation of this manuscript. The primary role of LLMs was to aid in polishing the text, which included improving grammar, refining sentence structure for clarity, and checking for stylistic consistency.
Appendix B Proofs
B.1 Proof of Theorem 2.10
Proof.
We begin with the definition of the number of realizations for a complete reasoning process.
Definition B.1.
$N_{\text{realize}}(\xi_{1},...,\xi_{L})=\prod_{l=1}^{L}|\mathcal{V}(\xi_{l},k_{l})|.$
Consider a single, fixed reasoning-level path $(\xi_{1},...,\xi_{L})$, where each step $\xi_{l}$ is realized by an average of $k$ tokens. By Assumption 2.9, the number of token-level realizations for this single path, $N_{\text{realize}}^{K}$, is bounded below:
$N_{\text{realize}}^{K}\geq\prod_{l=1}^{L}c^{k_{l}}=c^{\sum_{l=1}^{L}k_{l}}=c^{kL}=c^{K}.$ (1)
Next, we consider the total number of distinct, valid reasoning states, $|\mathcal{S}_{\text{reasoning}}^{(K)}|$. For complex problems, given a total token amount $K$, the valid reasoning paths can be realized in distinct lengths that are no longer than $K$, which means $|\mathcal{S}_{\text{reasoning}}^{(K)}|=\sum_{K^{{}^{\prime}}\leq K}N_{\text{realize}}^{K^{{}^{\prime}}}$. According to Equation (1), the number of realizations is bounded by an exponential term of the total number of tokens $K$, thus $|\mathcal{S}_{\text{reasoning}}^{(K)}|$ is of the same order with $N_{\text{realize}}^{K}$.
$|\mathcal{S}_{\text{reasoning}}^{(K)}|=\Theta(c^{K}).$ (2)
From Definition 2.8, the reasoning state density $\rho(K)$ is thus:
$\rho(K)=\frac{|\mathcal{S}_{\text{reasoning}}^{(K)}|}{|\mathcal{V}_{\text{semantic}}|}=\Theta(c^{K}).$ (3)
Assuming the semantic volume $\mathcal{V}_{\text{semantic}}$ for a given problem domain is fixed, the state density $\rho(K)$ thus grows exponentially with $K$.
Finally, we connect high density to the expected distance between adjacent points with the following lemma.
Lemma B.2.
(Nearest neighbor distance). Let points be distributed in a $d$-dimensional Euclidean space $\mathbb{R}^{d}$ according to a homogeneous Poisson point process with density $\rho$. The expected distance to the $k$-th nearest neighbor, denoted by $\mathbb{E}[R_{k}]$, is given by
$\mathbb{E}[R_{k}]=\frac{1}{(k-1)!}\left(\frac{1}{\rho C_{d}}\right)^{1/d}\Gamma\left(k+\frac{1}{d}\right),$
where $C_{d}=\frac{\pi^{d/2}}{\Gamma(\frac{d}{2}+1)}$ is the volume of a unit $d$-dimensional ball and $\Gamma(\cdot)$ is the Gamma function.
The proof is provided in Appendix B.2. Lemma B.2 implies the relation between the density and the expected distance of the points in a space. By only considering the nearest adjacent points, in a $D$-dimensional space, the expected distance $\mathbb{E}[\operatorname{dist}_{s_{i}\bowtie s_{j}}(s_{i},s_{j})]$ between adjacent nearest neighbor points $s_{i},s_{j}$ is inversely related to the density, approximately scaling as:
$\mathbb{E}_{s_{i}\bowtie s_{j}}[\operatorname{dist}(s_{i},s_{j})]\propto\rho^{-1/D}.$ (4)
Since $\rho(K)$ grows exponentially, it follows that:
$\lim_{K\to\infty}\mathbb{E}_{s_{i}\bowtie s_{j}}[\operatorname{dist}(s_{i},s_{j})]=0.$ (5)
When the expected distance between states approaches zero, it implies that for any state and any direction, another valid state can be found within an infinitesimally small neighborhood. This is the defining characteristic of a continuous space or manifold. Therefore, as the reasoning token amount $K$ increases, the discrete CoT-Space increasingly resembles a continuum, which validates the application of continuous optimization tools for its analysis.
This finishes the proof. ∎
B.2 Proof of Lemma B.2
Proof.
The proof is built upon Haenggi (2005) and  Miyagawa (2018).
First, we derive the probability density function, $f_{k}(r)$, for the distance to the $k$-th nearest neighbor. The cumulative distribution function (CDF), $F_{k}(r)$, is the probability that a $d$-dimensional ball $B_{r}$ of radius $r$ contains at least $k$ points. The volume of this ball is $V_{d}(r)=C_{d}r^{d}$.
The number of points $X$ in a region of volume $V$ follows a Poisson distribution with mean $\rho V$:
$P(X=x)=\frac{(\rho V)^{x}}{x!}e^{-\rho V}.$ (6)
The CDF is therefore $1$ minus the probability of finding fewer than $k$ points in the ball $B_{r}$:
$F_{k}(r)=P(\text{points in }B_{r}\geq k)=1-\sum_{x=0}^{k-1}P(X=x)=1-e^{-\rho C_{d}r^{d}}\sum_{x=0}^{k-1}\frac{(\rho C_{d}r^{d})^{x}}{x!}.$ (7)
The PDF is the derivative of the CDF with respect to $r$:
[…]
Using the product rule, the derivative simplifies due to a telescoping sum, yielding:
$f_{k}(r)=\frac{d(\rho C_{d})^{k}r^{dk-1}}{(k-1)!}e^{-\rho C_{d}r^{d}}.$ (9)
The expected value $\mathbb{E}[R_{k}]$ is defined as the integral of $r$ times its PDF over its domain:
[…]
Substituting the PDF we derived:
[…]
We combine the terms involving $r$ and move the constants outside the integral:
[…]
To solve the integral, we perform a change of variables. Let $t=\rho C_{d}r^{d}$. This implies:
[…]
The integral becomes:
[…]
The remaining integral is the definition of the Gamma function, $\Gamma(k+1/d)$. Thus:
[…]
Substituting this result back into the expression for $\mathbb{E}[R_{k}]$:
$\mathbb{E}[R_{k}]=\frac{d(\rho C_{d})^{k}}{(k-1)!}\left(\frac{\Gamma(k+1/d)}{d(\rho C_{d})^{k+1/d}}\right).$ (16)
The terms $d$ and $(\rho C_{d})^{k}$ cancel out, leaving:
$\mathbb{E}[R_{k}]=\frac{1}{(k-1)!}\frac{\Gamma(k+1/d)}{(\rho C_{d})^{1/d}}.$ (17)
Which can be written as the final result:
$\mathbb{E}[R_{k}]=\frac{1}{(k-1)!}\left(\frac{1}{\rho C_{d}}\right)^{1/d}\Gamma\left(k+\frac{1}{d}\right).$ (18)
This completes the proof. ∎
B.3 Proof of Theorem 2.12
Our goal is to show that both error components converge to zero as the reasoning state density $\rho(K)$ increases. Consider a $D$-dimensional semantic space. This analysis is predicated on a reasonable assumption about the locality of a single reasoning step.
Assumption B.3.
(Bounded step optimization). Any single reasoning step corresponds to a transition in the semantic space with an effective maximum radius $R_{\text{max}}$ and a minimum radius $R_{\text{min}}$. Thus, from a state $s$, the set of all possible subsequent states is contained within the region between two concentric $D$-dimensional hyperspheres centered at $s$ with radii $R_{\text{min}}$ and $R_{\text{max}}$. This region can be denoted as $B_{R}(s)$.
Following Assumption B.3, we continue to use the following definition to characterize the distribution structure of state points in the semantic space.
Definition B.4.
(Maximal void.) The maximal void $B_{R}^{\text{hole}}(s)$ is defined as the largest void hypersphere region within the bounded step radius. Formally, it is the largest hypersphere subspace of $B_{R}(s)$ that does not contain any state points:
$B_{R}^{hole}(s)=\underset{B}{\arg\max}\,\left\{r(B)\mid B\subseteq B_{R}(s)~\text{and}~B\cap\mathcal{S}=\emptyset\right\},$
where $r(\cdot)$ is the radius of a ball and $\mathcal{S}$ is the reasoning state space.
Let $R_{\text{hole}}$ be the radius of $B_{R}^{\text{hole}}(s)$. We further show the relationship between $R_{\text{hole}}$ and the reasoning state density $\rho$.
Lemma B.5.
(Maximal void radius and reasoning state density.) In a $D$-dimensional semantic space, the radius of the maximal void $R_{\text{hole}}$ and the density of reasoning state points $\rho(K)$ have the following relationship:
$R_{\text{hole}}=\mathcal{O}\left(\rho(K)^{-\frac{1}{D}}\right).$
Proof.
The proof relies on the connection between the average nearest-neighbor distance and the scale of the largest voids in a spatial point process.
As established in Theorem 2.10, which builds upon the theory of spatial point processes (Lemma B.2), the expected distance between adjacent reasoning states in a $D$-dimensional semantic space is inversely related to the state density $\rho(K)$:
$\mathbb{E}_{s_{i}\bowtie s_{j}}[\text{dist}(s_{i},s_{j})]\propto\rho(K)^{-1/D}.$ (19)
The radius of the maximal void, $R_{\text{hole}}$, represents the largest possible distance from a point within that void to the nearest available reasoning state. It can thus be understood as a measure of the worst-case nearest-neighbor distance in a local region.
In the study of stochastic geometry, for a homogeneous point process, the distribution of nearest-neighbor distances is concentrated around its mean. Consequently, the scaling of the maximal values is of the same order as the scaling of the expected value.
Given that the expected distance is proportional to $\rho(K)^{-1/D}$, it follows that the maximal void radius, representing the worst-case local distance, must adhere to the same scaling law. Therefore, we can conclude that:
$R_{\text{hole}}=\mathcal{O}\left(\rho(K)^{-\frac{1}{D}}\right).$ (20)
This completes the proof. ∎
With Lemma B.5, we then commence to analyze the convergence of the two continuum errors, respectively.
Analysis of the Angular Error $\epsilon_{A}(s)$
The angular error measures the difference in direction between discrete steps and continuous steps. Building upon the previous discussion, we can bound the angular error by the following result.
Lemma B.6.
(Convergence of the Angular Error). The expected angular error $\mathbb{E}\left[\epsilon_{A}(s)\right]$ is bounded above by a function that is inversely related to the reasoning state density $\rho(K)$:
$\mathbb{E}[\epsilon_{A}(s)]\leq\mathcal{O}\left(\rho(K)^{-\frac{1}{D}}\right).$
Proof.
The angular error $\epsilon_{A}(s)$ is defined as the angle between the ideal step vector $\vec{v}_{\text{ideal}}$ and the best achievable discrete step vector $\vec{v}_{\text{real}}$. This error is determined by the directional sparsity of available states in the semantic space.
The worst-case error occurs when the ideal vector $\vec{v}_{\text{ideal}}$ points directly towards the center of the maximal void, $B_{R}^{\text{hole}}(s)$, within the local neighborhood $B_{R}(s)$. Let the radius of this maximal void be $R_{\text{hole}}$. The ideal target state $p_{\text{ideal}}=s+\vec{v}_{\text{ideal}}$ is at the center of this void.
By definition, the best achievable subsequent state, $s_{j}^{*}$, must lie on or outside the boundary of this void. The vector difference $s_{j}^{*}-p_{\text{ideal}}$ represents the deviation from the ideal target.
Let’s consider the triangle formed by the points $s$ (current state), $p_{\text{ideal}}$ (ideal target), and $s_{j}^{*}$ (best achievable state). The angular error $\epsilon_{A}(s)$ is the angle at vertex $s$. Using the small-angle approximation for sine, which is valid as the space becomes dense, we can bound the angle:
$\epsilon_{A}(s)\approx\sin(\epsilon_{A}(s))\approx\frac{\|s_{j}^{*}-p_{\text{ideal}}\|}{\|\vec{v}_{\text{real}}\|}\leq\frac{\|s_{j}^{*}-p_{\text{ideal}}\|}{R_{\text{min}}}.$ (21)
In the worst-case scenario, the distance from the ideal point to the best real point $\|s_{j}^{*}-p_{\text{ideal}}\|$ is bounded by the properties of the void. This distance is on the order of the void’s radius, $R_{\text{hole}}$. Since the step size $\|\vec{v}_{\text{real}}\|$ is bounded and non-zero, the angular error is directly proportional to the radius of the maximal void:
$\epsilon_{A}(s)=\mathcal{O}(R_{\text{hole}}).$ (22)
Taking the expectation over all possible configurations of the state space, we have:
$\mathbb{E}[\epsilon_{A}(s)]=\mathcal{O}(\mathbb{E}[R_{\text{hole}}]).$ (23)
By combining this with Lemma B.5, we establish the bound on the expected angular error:
$\mathbb{E}[\epsilon_{A}(s)]\leq\mathcal{O}(\rho(K)^{-\frac{1}{D}}).$ (24)
As established in Theorem 2.10, the reasoning state density $\rho(K)$ grows exponentially with the total reasoning token amount $K$. Consequently, as $K\rightarrow\infty$, $\rho(K)\rightarrow\infty$, and thus $\lim_{K\rightarrow\infty}\mathbb{E}[\epsilon_{A}(s)]=0$. This demonstrates that the direction of the best discrete step converges to the direction of the ideal continuous gradient.
This completes the proof.
Analysis of the Magnitude Error $\epsilon_{M}(s)$
The magnitude error quantifies how well the length of the discrete step matches the length of the ideal step. This insight provides a direct way to bound this error.
Lemma B.7.
(Convergence of the Magnitude Error). The expected magnitude error $\mathbb{E}\left[\epsilon_{M}(s)\right]$ is also bounded above by a function that is inversely related to the reasoning state density $\rho(K)$:
$\mathbb{E}[\epsilon_{M}(s)]\leq\mathcal{O}\left(\rho(K)^{-\frac{1}{D}}\right).$
Proof.
By the reverse triangle inequality, the magnitude error is bounded by the norm of the vector difference:
$\epsilon_{M}(s)=\left|\|\vec{v}_{\text{real}}\|-\|\vec{v}_{\text{ideal}}\|\right|\leq\|\vec{v}_{\text{real}}-\vec{v}_{\text{ideal}}\|.$ (25)
[…]
In the worst-case scenario, the ideal target $p_{\text{ideal }}$ falls within the largest void. The closest available state $s_{j}^{*}$ must lie on or outside the boundary of this void. The distance between any point inside a hypersphere and its boundary is at most its diameter, $2R_{\text{hole}}$. Therefore, the distance between the ideal target and the best real state is bounded:
$\|s_{j}^{*}-p_{\text{ideal}}\|\leq 2R_{\text{hole}}.$ (26)
Combining these steps, we get a bound for the magnitude error:
$\epsilon_{M}(s)\leq 2R_{\text{hole}}.$ (27)
Taking the expectation and applying Lemma B.5, we find that the expected magnitude error is bounded by the same scaling law as the angular error:
$\mathbb{E}[\epsilon_{M}(s)]\leq 2\mathbb{E}[R_{\text{hole}}]=\mathcal{O}\left(\rho(K)^{-1/D}\right).$ (28)
As $K\rightarrow\infty$, $\rho(K)\rightarrow\infty$, and thus $\lim_{K\rightarrow\infty}\mathbb{E}\left[\epsilon_{M}(s)\right]=0$.
This completes the proof. ∎
B.4 Proof of Theorem 3.1
Proof.